# Video Frame Interpolation with Region-Distinguishable Priors from SAM

### v1

@@ 3.2 Region-Distinguishable Priors(RDPs)
Figure 3: The Overview of RHFFM, which first exploits DRPs to enhance image features via DRPFN (Eq. 6), and then refine it via refinement (Eq. 7). ${f_{i,l},s_{i,l}}$ are the image feature and DRP feature of the $i$-th frame of the $l$-th layer, respectively.
3.3 HRFFM
As indicated in Sec. 3.1, standard motion-based VFI conducts multi-scale feature extraction before motion estimation. Thus, we put the obtained RDPs (from Sec. 3.2) into each layer of image feature extraction as shown in Fig. 2. The fusion consists of three stages, including RDP feature extraction, RDP-guided Feature Normalization (RDPFN), and RDP residual learning, as exhibited in Fig. 3.
@@ 3.3 HRFFM
where $\mathcal{R}_{l}$ is the RDPFN operation in the $l$-th layer. The details of RDPFN will be introduced in Sec. 3.4.
Figure 4: The Overview of RDPFN. It utilizes both RDP features and image features as inputs. It employs a combination of long- and short-range operations to extract impactful features, facilitating the prediction of region-aware normalization parameters. This approach ensures that features within the same instance exhibit similarity, thereby enhancing the effect of subsequent modules.
Moreover, we recognize that segmentation results obtained from SAM may contain errors when dealing with diverse real-world images. Consequently, additional refinement operations are essential to enhance the features derived from RDPFN, rendering them more adaptable for subsequent motion estimation and frame synthesis. In our study, we have identified a refinement operation that enhances robustness and is accomplished through a spatial-channel convolution fusion in a residual manner, as
@@
3.4 RDP-guided Feature Normalization
@@
4 Experiments
Methods Vimeo90K UCF101 SNU-FILM parameters runtime
Easy Medium Hard Extreme (millions) (seconds)
DQBC [52] 36.57/ 0.9817 35.44/ 0.9700 40.31/0.9909 36.25/ 0.9799 30.94/ 0.9378 25.61/ 0.8648 18.3 0.206
IFRNet [21] 36.20/0.9808 35.42/0.9698 40.10/0.9906 36.12/0.9797 30.63/0.9368 25.27/0.8609 19.7 0.79
EBME [17] 36.19/0.9810 35.41/0.9700 40.28/0.9910 36.07/ 0.9800 30.64/0.9370 25.40/0.8630 3.9 0.08
ABME [41] 36.18/0.9805 35.38/0.9698 39.59/0.9901 35.77/0.9789 30.58/0.9364 25.42/0.8639 18.1 0.22
SoftSplat [34] 36.10/0.9700 35.39/0.9520
VFIformer [29] 36.38/0.9811 35.34/0.9697 40.16/0.9907 35.92/0.9793 30.20/0.9337 24.80/0.8551 24.17 0.63
@@ 4 Experiments
UPR-Net [18] 36.02/0.9800 35.40/0.9698 40.40/ 0.9910 36.15/0.9797 30.70/0.9364 25.53/0.8631 1.65 0.05
@@
Table 1: Qualitative (PSNR/SSIM) comparisons between VFI baselines and their implementation with our strategy (${}_{ours}$) on UCF101 [46], Vimeo90K [49], and SNU-FILM [8] benchmarks. The best result and the second best are boldfaced and underlined, respectively. Our strategy enhances the performance of various representative VFI methods. When combined with our approach, these methods can surpass current SOTA approaches. Moreover, note that our approach does not introduce a significant increase in computation cost. For running time, we test all models under 640 × 480 resolution, and average the running time by 100 iterations.
4.1 Datasets
Training dataset. The Vimeo90K dataset [49] contains 51,312 triplets with a resolution of 448$\times$256 for training. We augment the training images by randomly cropping 256$\times$256 patches. We also apply random flipping, rotating, and reversing the order of the triplets for augmentation.
  • UCF101 [46]: The test set of UCF101 contains 379 triplets with a resolution of 256$\times$256. UCF101 contains a large variety of human actions.
  • Vimeo90K [49]: The test set of Vimeo90K contains 3,782 triplets with a resolution of 448$\times$256.
  • SNU-FILM [8]: This dataset contains 1,240 triplets, and most of them are of a resolution of around 1280$\times$720. It contains four subsets with increasing motion scales – easy, medium, hard, and extreme.
4.2 Implementation Results
We evaluate our proposed HRFFM with RDPs to enhance the performance of current representative VFI baselines, including VFIformer [29], UPR-Net [18] and M2M-PWC [12]. To ensure a fair comparison, we report results by implementing the officially released source code and training models under unified conditions on the same machine, rather than replicating results from the original papers. We maintain the original model architecture and loss function, incorporating our method into the feature encoder, as illustrated in Fig. 2.
@@ 4.2 Implementation Results
Figure 5: Visual comparison on SNU-FILM [8]. Three rows, from top to bottom, represent the comparison results for VFIformer, UPR-Net, and M2M-PWC. The highlighted boxes indicate positions where our model demonstrates superior performance.
4.3 Comparison with VFI Baselines
@@
Quantitative comparison. The comparison results are presented in Tab. 1, where we integrate our proposed approach with VFI baselines to assess performance improvements. It is observed that almost all baselines exhibit enhanced results across all testing sets when our strategy is applied, with only a minimal increase in parameters and computation costs. Notably, our method demonstrates a substantial improvement of 0.31dB on Vimeo90K for the robust baseline, VFIformer. Additionally, for other methods, there is an improvement of more than 0.1dB, which is significant in the context of VFI tasks where performance has almost approached the upper limit.
@@
Moreover, we conduct a comparison of our model with several other SOTA VFI models, including DQBC [52], IFRNet [21], EBME [17], ABME [41], and SoftSplat [34], as outlined in Tab. 1. The results reveal that when integrated with our strategy, the chosen VFI baselines can outperform these competitive SOTA approaches.
@@
Figure 6: The comparison between baselines and ours in terms of the downstream video segmentation task. The top row is three continuous frames for visualization. The second and third rows are the segmentation results on the input and synthesized intermediate frames with the baseline and ours, respectively.
[…]
Settings Vimeo90K SNU-FILM
@@
Ours with O.H. 35.52 40.14/0.9908 35.86/0.9791 30.46/0.9354 25.38/0.8619
Ours with L.E. 35.40 40.08/0.9908 35.79/0.9790 30.34/0.9349 25.26/0.8610
Ours w/o S.O. 35.52 40.17/0.9908 35.86/0.9791 30.46/0.9353 25.38/0.8617
Ours w/o R.L. 35.54 40.12/0.9908 35.88/0.9791 30.51/0.9357 25.42/0.8629
Ours with CNN 35.39 40.03/0.9907 35.79/0.9788 30.41/0.9345 25.42/0.8615
Ours with Trans. 35.53 40.15/0.9908 35.85/0.9790 30.46/0.9351 25.35/0.8615
Full 35.57 40.15/0.9908 35.89/0.9791 30.48/0.9354 25.38/0.8619
Table 2: Ablation study results for the proposed strategy.
4.4 Ablation Study
In this section, we perform various ablation studies to examine different components in our proposed method. All ablation tests are carried out using UPR-Net, and we present qualitative results from training for 100,000 iterations.
@@ 4.4 Ablation Study
Effect of Mixture Gaussian embedding. Mixture Gaussian embedding serves as a crucial representation for distinguishing objects between two frames, playing a pivotal role in adapting SAM outputs for an arbitrary number of instances. To investigate the impact of Mixture Gaussian embedding, we replaced it with alternative methods, including naive one-hot encoding or learnable embeddings. Both alternatives require assuming a maximum instance number, denoted as “Ours with O.H.” and “Ours with L.E.”, respectively. The results, presented in Tab. 2, indicate that their performance is lower than the results achieved with Mixture Gaussian embedding, highlighting the effect of the proposed approach outlined in Sec. 3.2.
@@
Effect of softmax operation and residual learning in HRFFM. After the feature extraction for RDP in each layer, the softmax operation ensures the consistency of feature representations at different scales. Additionally, to mitigate the impact of SAM errors on subsequent feature fusion, a residual learning component is incorporated after RDPFN. To assess their effectiveness, we trained two models without the softmax operation and residual learning, labeled as “Ours w/o S.O.” and “Ours w/o R.L.”, respectively. As depicted in Tab. 2, the performance of both models is lower than the original full setting, underscoring the rationality of the softmax operation and residual learning in HRFFM.
Figure 7: Visual comparisons of ablation studies on Vimeo90K [49]. Six rows, from top to bottom, represent the comparison results for “Ours with O.H.”, “Ours with L.E.”, “Ours w/o S.O.”, “Ours w/o R.L.”, “Ours with CNN”, and “Ours with Trans.”.
@@
Effect of parallel CNN and transformer blocks in RDPFN. RDPFN is designed to leverage both long- and short-range dependencies, formulating normalization parameters for regions with varying shapes and areas. To demonstrate the effectiveness of this parallel setting, we trained two models with only a convolutional layer and a Transformer layer in RDPFN, labeled as “Ours with CNN” and “Ours with Trans.”, respectively. The results in Tab. 2 indicate that removing either component leads to an overall performance degradation, underscoring the necessity of the parallel CNN and Transformer strategy in formulating suitable region-aware normalization parameters.
In addition to quantitative comparisons, we also present visual comparisons. As shown in Fig. 7, the intermediate frames generated by six ablation studies and our method are shown in the last two columns. Obviously, our method produces better results than the others.
4.5 User Study

### v2

@@ III-B Region-Distinguishable Priors (RDPs)
Note that the codebook entries are pre-defined and not learned. Their primary role is to provide distinct values to represent different objects, which are used to generate the RDPs. The learning process is carried out by the subsequent HRFFM network, which learns to effectively utilize these RDPs through feature fusion. Further details are provided in the next section.
Figure 3: The Overview of HRFFM, which first exploits RDPs to enhance image features via RDPFN (Eq. 6), and then refine it via refinement (Eq. 7). $f_{i,l},s_{i,l}$} are the image feature and RDP feature of the $i$-th frame of the $l$-th layer, respectively. $\oplus$ means concatenating.
III-C HRFFM
As indicated in Sec. III-A, standard motion-based VFI conducts multi-scale feature extraction before motion estimation. Thus, we put the obtained RDPs (from Sec. III-B) into each layer of image feature extraction as shown in Fig. 2. The fusion consists of three stages, including RDP feature extraction, RDP-guided Feature Normalization (RDPFN), and RDP residual learning, as exhibited in Fig. 3.
@@ III-C HRFFM
where $\mathcal{R}_{l}$ is the RDPFN operation in the $l$-th layer. The details of RDPFN will be introduced in Sec. III-D.
Figure 4: The Overview of RDPFN. It utilizes both RDP features and image features as inputs. It employs a combination of long- and short-range operations to extract impactful features, facilitating the prediction of region-aware normalization parameters. This approach ensures that features within the same instance exhibit similarity, thereby enhancing the effect of subsequent modules. $\oplus$} means concatenating , $\otimes$} means dot producting.
Moreover, we recognize that segmentation results obtained from SAM2 may contain errors when dealing with diverse real-world images. Consequently, additional refinement operations are essential to enhance the features derived from RDPFN, rendering them more adaptable for subsequent motion estimation and frame synthesis. In our study, we have identified a refinement operation that enhances robustness and is accomplished through a spatial-channel convolution fusion in a residual manner, as
@@
III-D RDP-guided Feature Normalization
@@
III-E Details of Our Framework’s Components
Our framework comprises two key components: 1) Region-Distinguishable Priors (RDPs), represented as spatially-varying Gaussian mixture models, and 2) the core Hierarchical Region-aware Feature Fusion Module (HRFFM), enhancing feature representations for accurate motion estimation.
RDPs come out of the generated mask $M_{i}$ from SAM, and a Gaussian distributed sampler $\mathcal{G}$, written as ${S}_{i}=\mathcal{G}(M_{i})$, where ${S_{i}}\in\mathbb{R}^{H\times W\times c}$. In our experiments, $c$ is 3, the same dimension as the input frame. An ablation study to analyze the impact of this channel number will be given in Sec. IV-D, verifying that 3 is the optimal hyper-parameter.
The HRFFM consists of three parts, including RDP feature extraction, RDP-guided Feature Normalization (RDPFN), and RDP residual learning. In the procedure of RDP feature extraction, two multi-scale encoders are set for RDPs and input frames, respectively. They both have a series of convolution layers, guaranteeing the same size of frame/RDP encoder’s outputs. HRFFM can be adopted as a play-and-plug module for different VFI frameworks, and we take UPR-Net as an example here. As shown in Tab. I, the feature encoder of UPR-Net has three convolutional stages: stage-0, stage-1, and stage-2. Each stage consists of four convolution layers, and the first layers in stage-1 and stage-2 perform down-sampling. Analogously, the RDP feature encoder, UPR-Netours, has three convolution stages. Each stage consists of two convolution layers as shown in Tab. II, and the first layers of stage-1 and stage-2 perform down-sampling. One difference is that the RDP feature extraction discards activation functions to speed up training.
In RDPFN, the $l$-th layer feature of frame $f_{i,l}$ is first transformed to $\bar{f}_{i,l}$ by a batch normalization operation. Then we concatenate it with the $l$-th layer feature of RDP $s^{\prime}_{i,l}$ to get $c_{i,l}$-D embeddings (where $c_{i,l}=c_{s^{\prime}_{i,l}}+c_{\bar{f}_{i,l}}$, $c_{s^{\prime}_{i,l}}$ and $c_{\bar{f}_{i,l}}$ represent the feature channels of $s^{\prime}_{i,l}$ and $\bar{f}_{i,l}$, respectively), and send it to: (1) the CNN block $\mathcal{K}_{l}$ that consists of two convolution layers (the strides are both set as 1) with output channel number as $c_{i,l}$; (2) the Transformer block $\mathcal{T}_{l}$ whose procedure is written as:
[…]
where $x_{i,l}$ is the input feature of $\mathcal{T}_{l}$, “Atten” is the multi-head self-attention (the number of heads is 2), and “FeedForward” is the feed-forward module in the transformer. Moreover, the learnable fusion $\mathcal{A}_{l}$, normalization prediction networks $\mathcal{B}_{\alpha}$ and $\mathcal{B}_{\beta}$ are both convolution layers with the stride of 1. The output of RDPFN $\hat{f}_{i,l}$ concatenates with $f_{i,l}$ and goes through a spatial channel convolution fusion to achieve the final augmented frame features $f^{\prime}_{i,l}$.
TABLE I: Feature encoder of input frames in UPR-Net [Alpher39].
Layer Type Activation Kernel Stride Padding Output Size
stage-0 Input Fearure - - - - $H\times W\times 3$
Convolution*4 LeakyReLu 3 1 1 $H\times W\times 16$
stage-1 Convolution LeakyReLu 3 2 1 $\frac{H}{2}\times\frac{W}{2}\times 32$
Convolution*3 LeakyReLu 3 1 1 $\frac{H}{2}\times\frac{W}{2}\times 32$
stage-2 Convolution LeakyReLu 3 2 1 $\frac{H}{4}\times\frac{W}{4}\times 64$
Convolution*3 LeakyReLu 3 1 1 $\frac{H}{4}\times\frac{W}{4}\times 64$
TABLE II: The feature encoder of RDPs in UPR-Netours.
Layer Type Activation Kernel Stride Padding Output Size
stage-0 RDPs Fearure - - - - $H\times W\times 3$
Convolution*2 - 3 1 1 $H\times W\times 16$
stage-1 Convolution - 3 2 1 $\frac{H}{2}\times\frac{W}{2}\times 32$
Convolution - 3 1 1 $\frac{H}{2}\times\frac{W}{2}\times 32$
stage-2 Convolution - 3 2 1 $\frac{H}{4}\times\frac{W}{4}\times 64$
Convolution - 3 1 1 $\frac{H}{4}\times\frac{W}{4}\times 64$
@@ III-E Details of Our Framework’s Components
IV Experiments
IV-A Datasets
@@
  • UCF101 [Alpher42]: The test set of UCF101 contains 379 triplets with a resolution of $\times256$. UCF101 contains a large variety of human actions.
@@
[…]
@@
Figure 5: Visual comparison on SNU-FILM [ChannelM2020]. Three rows, from top to bottom, represent the comparison results for VFIformer, UPR-Net, and M2M-PWC. The highlighted boxes indicate positions where our model demonstrates superior performance.
TABLE III: Quantitative (PSNR/SSIM) comparisons between VFI baselines and their implementation with our strategy (ours) on Vimeo90K [Alpher41], and the “extreme” subset of SNU-FILM [ChannelM2020] benchmarks. Our strategy enhances the performance of various representative VFI methods. Note that our approach does not introduce a significant increase in computation cost. For running time, we test all models under 640 × 480 resolution, and average the running time by 100 iterations. Moreover, the results on Vimeo90K were obtained by training solely on the Vimeo90K dataset, while the SNU-FILM Extreme results were achieved by further fine-tuning the model on the training set augmented with the Easy, Medium, and Hard subsets, highlighting its improvement in challenging scenarios.
Methods Vimeo90K SNU-FILM parameters runtime
Extreme (millions) (seconds)
UPR-Net [Alpher39] 36.02/0.9800 25.67/0.8697 1.65 0.05
UPR-Netours 36.19/0.9806 25.94/0.8749 2.64 0.13
M2M-PWC [Alpher40] 35.27/0.9771 25.26/0.8663 7.61 0.04
M2M-PWCours 35.37/0.9775 25.47/0.8682 10.65 0.04
AMT [li2023amt] 36.50/0.9816 25.72/ 0.8723 29.22 0.06
AMTours 36.61/0.9828 25.87/0.8772 36.74 0.14
Bim-VFI [seo2025bim] 35.12/0.9753 24.63/0.8551 6.56 0.05
Bim-VFIours 35.31/0.9763 24.75/0.8573 10.82 0.06
TABLE IV: Quantitative (PSNR/SSIM) comparisons between VFI baselines and VFIformer’s implementation with our strategy (ours) on Vimeo90K [Alpher41] . The best result is boldfaced . When combined with our approach, this method (VFIformer) can surpass current SOTA approaches.
VFIformer VFIformerours EMA-VFI [zhang2023extracting] DQBC  [dqbc2023] IFRNet  [ifrnet2022] EBME  [ebme2022] ABME  [abme2021] SoftSplat  [softsplat2020] SGM-VFI [liu2024sparse] TTVFI  [liu2023ttvfi] InterpAny  [zhong2024clearer]
36.38 36.69 36.64 36.57 36.20 36.19 36.18 36.10 32.50 36.54 35.51
0.9811 0.9826 0.9819 0.9817 0.9808 0.9810 0.9805 0.9700 0.9689 0.9819 0.9779
IV-B Implementation Results
We evaluate our proposed HRFFM with RDPs to enhance the performance of current representative VFI baselines, including VFIformer [Alpher19], UPR-Net [Alpher39], M2M-PWC [Alpher40], AMT [li2023amt], and Bim-VFI [seo2025bim]. To ensure a fair comparison, we report results by implementing the officially released source code and training models under unified conditions on the same machine, rather than replicating results from the original papers. We maintain the original model architecture and loss function, incorporating our method into the feature encoder, as illustrated in Fig. 2.
IV-C Comparison with VFI Baselines
@@
Quantitative comparison. Comparison results are presented in Tab. III, where we integrate our proposed approach with VFI baselines to assess performance improvements. It is observed that almost all baselines exhibit enhanced results across all testing sets when our strategy is applied, with only a minimal increase in parameters and computation costs. Notably, our method demonstrates a substantial improvement of 0.31dB on Vimeo90K for the robust baseline, VFIformer (presented in Tab. IV). For other methods, there is an improvement of more than 0.1dB, which is significant in the context of VFI tasks where performance has almost approached the upper limit.
@@
Moreover, we conduct a comparison with several current SOTA VFI methods, including EMA-VFI [zhang2023extracting], DQBC [dqbc2023], IFRNet [ifrnet2022], EBME [ebme2022], ABME [abme2021], SoftSplat [softsplat2020], SGM-VFI [liu2024sparse], TTVFI [liu2023ttvfi], and InterpAny [zhong2024clearer], as outlined in Tab. IV. The results reveal that when integrated with our strategy, the chosen VFI baseline, i.e., VFIformer, can outperform these competitive SOTA approaches.
Furthermore, we conduct experiments on the large-scale benchmark LAVIB [stergiou2024lavib]. Specifically, we use LAVIB to evaluate our method alongside several representative VFI approaches. In addition to standard metrics, we also include FVD [unterthiner2019fvd], which measures the distance between generated outputs and ground truth across multiple dimensions. In particular, we randomly sample a sufficient number of training and evaluation examples from the corresponding datasets, ensuring that the train-test split follows the official setting and remains consistent across all competing methods. As shown in Tab. V, our method continues to outperform others on this dataset.
TABLE V: The comparison between VFI baselines and their counterparts enhanced with our method, evaluated across more datasets and additional metrics.
Methods Vimeo90K SNU-FILMextreme LAVIB
[…]
UPR-Net 99.9108 179.31 31.86/0.9420
UPR-Netours 97.3529 176.85 31.97/0.9563
@@
Moreover, even in the presence of occlusions (as shown in the foot region of Fig. 7), our approach can effectively distinguish the corresponding regions via RDPs obtained from SAM2, preserving clear boundaries and reducing interpolation errors in those areas.
@@
Figure 6: Comparisons between baselines and ours in terms of the downstream video segmentation task. The top row is three continuous frames for visualization. The second and third rows are the segmentation results on the input and synthesized intermediate frames with the baseline and ours, respectively.
Evaluation with downstream tasks. The VFI capability can be leveraged for various downstream tasks, including video segmentation. Large temporal gaps in videos can disrupt the effective propagation of semantic information. To assess the effect of our framework in terms of its impact on downstream video segmentation tasks, we employ the SOTA video segmentation approach SAM-Track[cheng2023segment]. The results, presented in Fig. 6, showcase three consecutive frames in the first row, with segmentation results of synthesized intermediate frames generated by VFIformer and VFIformerours in the second and third row, respectively. It is evident that the intermediate frames produced by our model exhibit more accurate segmentation. Our method’s results enhance better temporal propagation among frames and can even rectify incorrect segmentation results in the first frame. For instance, the dog in the second row is not clearly separated from the shadow on the ground, whereas in the third row, the separation is more distinct.
Figure 7: The t-SNE visualization comparison about the intermediate features between the baseline VFI and our enhanced version, VFIours. Two colors represent the feature distributions of the same object in two different frames. As shown, the left t-SNE figure (c) exhibits more isolated regions, indicating that the two feature distributions are more separated. In contrast, the right t-SNE figure (d) shows a greater overlap between the distributions, demonstrating improved feature consistency across frames. Moreover, as shown in the bottommost figure (e) and (f), we select two pixels within the red bounding box for feature comparison. The two dark-colored points represent the features of this pixel in the two consecutive frames synthesized by the VFI model, whereas the yellow and green points correspond to those generated by our method. Clearly, the feature distance between the two frames produced by our approach is substantially smaller.
Feature-level visualization for alignment analysis. We propose a feature-level visualization strategy to illustrate the effectiveness of our method in achieving motion alignment and thus enhancing interpolation quality. To create this visualization, we first extract feature maps from the start and end frames and identify feature regions belonging to the same object, referred to as “feature region A” and “feature region B”. We then apply the t-SNE technique to project all feature vectors from regions A and B into a 2D space. If alignment is successful, these feature vectors—representing the same object—should be closely aligned and overlapped. As shown in Fig. 7, feature vectors extracted by our encoder from regions A and B (the leg area within the red box in im0 and im1) exhibit stronger alignment compared to the baseline. For clearer analysis, we select the feature of a single pixel within the red bounding box for comparison. As shown in the visualization, the distance between these two features produced by our method is significantly smaller, indicating that our approach effectively enhances feature alignment. Additionally, we compute the overlap mIoU between these two groups of feature vectors in the 2D space, with values of 32% (ours) and 15% (baseline), further confirming that our method achieves better feature alignment for the same object across different frames.
Moreover, we visualize the optical flow obtained from the deep features with a few layers. As shown in Fig. 1, different regions exhibit distinct flow values with clear boundaries between them. This indicates that the regions have different feature representations, as reflected by the variations in their decoded flow values. This separation could facilitate interpolation by preventing incorrect alignment across varying objects.
Comparison with generative-model-based methods. As indicated in the introduction, our method differs in its objectives from generative-model-based approaches, which often fall short in achieving high fidelity. To support this claim, we provide the evaluation results of a SOTA and representative generative frame interpolation method, FCVG [zhu2025generative]. Its performance on various datasets is as follows: 25.52/0.8554 on Vimeo90K, 18.35/0.6876 on UCF101, 24.29/0.8021 on SNU-FILM (Easy), 21.85/0.7516 on SNU-FILM (Medium), 19.89/0.7237 on SNU-FILM (Hard), and 18.44/0.6618 on SNU-FILM (Extreme). These results are clearly lower than those of restoration-based methods, as shown in Tab. III. Thus, while generative methods are capable of producing photo-realistic frames, their fidelity to the ground truth remains relatively low.
IV-D Ablation Study
In this section, we perform various ablation studies to examine different components in our proposed method. The first ablation test ( the corresponding results in Tab. VIII) was trained to completion, while all subsequent ablation experiments were truncated at 100,000 iterations.
Effect of HRFFM. We have designed a comprehensive framework that leverages semantic segmentation features derived from SAM2, specifically RDPs, to distinguish objects in images. Additionally, we meticulously crafted a module that fuses image and segmentation features to enhance the accuracy of flow estimation, thereby improving the performance of intermediate frame interpolation. In this section, we explore two crucial questions: (1) whether RDPs are indispensable components of the proposed framework, and (2) whether the performance gains of HRFFM are solely due to the increased network depth or can be attributed to the specific design of the framework. To address these questions, we conducted two experiments. First, we removed the RDPs from the framework, as illustrated in $s^{\prime}_{i,l}$ of Fig. 3. Second, we replaced our HRFFM with a simple two-layer convolutional neural network while keeping all other conditions unchanged. The results of these experiments are presented in Tab. VIII, where UPR-NetNoRDPs and UPR-Net2conv correspond to these two experimental settings, respectively. The results clearly show that the model incorporating segmentation features outperforms the model relying solely on image features, as demonstrated by the comparative performance of UPR-Netours and UPR-NetNoRDPs. Furthermore, our proposed framework exhibits a substantial performance advantage over the simplistic two-layer convolutional neural network, indicating that the observed improvement is not merely a consequence of increased network depth, but rather the result of the synergistic integration of image and segmentation features.
TABLE VI: Ablation study results. We calculate PSNR, LPIPS, and NIQE values of different ablation settings in the Vimeo90K dataset. boldfaced indicates the best results.
Settings PSNR LPIPS NIQE
Ours with O.H. 35.52 0.023 26.95
Ours with L.E. 35.40 0.0233 27.15
Ours w/o S.O. 35.52 0.0152 24.09
Ours w/o R.L. 35.54 0.0154 24.04
Ours with CNN 35.39 0.0153 24.19
Ours with Trans. 35.53 0.0154 24.00
Full 35.57 0.0148 23.88
TABLE VII: Ablation study results for the proposed strategy. We calculate LPIPS values of different ablation settings in the challenging high-resolution SNU-FILM dataset.
Settings Easy Medium Hard Extreme
Ours with O.H. 0.0191 0.0337 0.0633 0.1178
Ours with L.E. 0.0195 0.034 0.0637 0.1159
Ours w/o S.O. 0.012 0.0218 0.0444 0.0898
Ours w/o R.L. 0.0123 0.0222 0.0452 0.0911
Ours with CNN 0.012 0.0218 0.0446 0.0901
Ours with Trans. 0.0117 0.0216 0.0447 0.0906
Full 0.0111 0.0205 0.0426 0.0877
TABLE VIII: Quantitative (PSNR/SSIM) comparisons between UPR-Net , UPR-Netours ,UPR-NetNoRDPs and UPR-Net2conv on Vimeo90K, UCF101, and SNU-FILM.
Methods Vimeo90K UCF101 SNU-FILM
@@ IV-D Ablation Study
UPR-Net [Alpher39] 36.02/0.9800 35.40/0.9698 40.40/0.9910 36.15/0.9797 30.70/0.9364 25.53/0.8631
UPR-Netours 36.19/0.9806 35.45/0.9699 40.43/0.9911 36.19/0.9798 30.80/0.9370 25.64/0.8643
UPR-NetNoRDPs 36.18/0.9806 35.44/0.9699 40.37/0.9910 36.15/0.9797 30.73/0.9370 25.57/0.8640
UPR-Net2conv 36.08/0.9803 35.41/0.9699 34.02/0.9745 32.81/0.9661 29.31/0.9279 24.99/0.8564
@@
Effect of Mixture Gaussian embedding. Mixture Gaussian embedding serves as a crucial representation for distinguishing objects between two frames, playing a pivotal role in adapting SAM2 outputs for an arbitrary number of instances. To investigate the impact of Mixture Gaussian embedding, we replaced it with alternative methods, including naive one-hot encoding or learnable embeddings. Both alternatives require assuming a maximum instance number, denoted as “Ours with O.H.” and “Ours with L.E.”, respectively. The results, presented in Tab. VI, indicate that their performance is lower than the results achieved with Mixture Gaussian embedding, highlighting the effect of the proposed approach outlined in Sec. III-B.
Figure 8: Visual comparisons of ablation studies on Vimeo90K [Alpher41].
@@
Effect of softmax operation and residual learning in HRFFM. After the feature extraction for RDP in each layer, the softmax operation ensures the consistency of feature representations at different scales. Additionally, to mitigate the impact of SAM2 errors on subsequent feature fusion, a residual learning component is incorporated after RDPFN. To assess their effectiveness, we trained two models without the softmax operation and residual learning, labeled as “Ours w/o S.O.” and “Ours w/o R.L.”, respectively. As depicted in Tab. VI, the performance of both models is lower than the original full setting, underscoring the rationality of the softmax operation and residual learning in HRFFM.
TABLE IX: The ablation study results for the Gaussian mixture channel choice. Different Gaussian mixture channels’ results on three datasets, Vimeo90K[Alpher41], UCF101 [Alpher42], and Middlebury[middlebury], are given.
Gaussian mixture channels Vimeo90K UCF101 Middlebury
3 35.57/0.9783 35.39/0.9696 37.49/0.9854
24 35.45/0.9780 35.31/0.9695 37.32/0.9847
96 35.39/0.9777 35.32/0.9695 37.28/0.9848
@@
Effect of parallel CNN and transformer blocks in RDPFN. RDPFN is designed to leverage both long- and short-range dependencies, formulating normalization parameters for regions with varying shapes and areas. To demonstrate the effectiveness of this parallel setting, we trained two models with only a convolutional layer and a Transformer layer in RDPFN, labeled as “Ours with CNN” and “Ours with Trans.”, respectively. The results in Tab. VI indicate that removing either component leads to an overall performance degradation, underscoring the necessity of the parallel CNN and Transformer strategy in formulating suitable region-aware normalization parameters. To further validate the conclusions from the ablation experiments, we conduct additional ablation experiments on the SNU-FILM datasets (shown in Tab. VII). These results also demonstrate that our proposed strategy outperforms the ablated settings. In addition to quantitative comparisons, we also present visual comparisons. As shown in Fig. 8, the intermediate frames generated by six ablation studies and our method are shown in the last two columns. Obviously, our method produces better results than the others.
TABLE X: Quantitative (PSNR/SSIM) comparisons between UPR-Net and UPR-Netours on Vimeo90K, UCF101, and SNU-FILM, whose data is degraded by the low-light.
Methods Vimeo90K UCF101 SNU-FILM
Easy Medium Hard Extreme
UPR-Net 42.06/0.9870 40.15/0.9759 44.49/0.9937 40.35/0.9850 34.97/0.9503 29.81/0.8895
UPR-Netours 42.19/0.9877 40.38/0.9763 44.64/0.9938 40.51/0.9852 35.17/0.9512 30.07/0.8910
The influence of Gaussian mixture channels. As described in Sec. III-E, the RDPs are represented as 3D Gaussian mixtures. We conducted an ablation study to explore the impact of different Gaussian mixture dimensions. The experiments were carried out using the UPR-Net framework, with each ablation setting trained for 100,000 iterations. The results, presented in Tab. IX, indicate that across all three datasets, when the Gaussian mixture dimension is set to 24 or 96, the interpolation performance is not as effective as with our original setting of 3. This confirms that increasing the Gaussian mixture dimension does not necessarily yield better results. We think there are two main reasons for this phenomenon. First, increasing the dimension raises model complexity, resulting in more parameters to optimize, which is challenging given the fixed number of training samples across experiments. Second, higher dimensionality introduces more randomness in the RDPs, making the learning process harder due to an increased degree of ill-posedness. In summary, increasing the dimension intensifies the learning difficulty, causing poor alignment and consequently incorrect motion interpolation results.
TABLE XI: Comparison of our method using SAM2 and alternative segmentation models, i.e., including the original SAM2 (UPR-Netours), a semantic segmentation variant (UPR-Netsemantic), and a lightweight version, Tiny-SAM (UPR-Nettinysam).
Methods Vimeo90K UCF101
$\uparrow PSNR\uparrow$
UPR-Net [Alpher39] 35.49 35.27
UPR-Netours 35.78 35.39
UPR-Netsemantic 35.65 35.30
UPR-Nettinysam 35.77 35.39
The results with efficient SAM2. The primary goal of this paper is to demonstrate the effectiveness of using RDPs from SAM2 to enhance the performance of existing VFI methods—an approach that has not been explored or implemented in prior work. Meanwhile, we also note that inference efficiency is crucial for real-world applications. Notably, recent advancements in SAM, such as Tiny-SAM [shu2025tinysam], have significantly improved inference speed. Therefore, we have conducted an additional experiment using Tiny-SAM to obtain our RDPs. The results, presented in Tab. XI, show that our method with Tiny-SAM still improves baseline performance, highlighting its potential for practical use. Note that due to limited time and resources, all experiments in Tab. XI were truncated at 200,000 iterations.
The results with semantic segmentation priors. We think that semantic information alone may not be effective for interpolation, as objects sharing the same semantic label can exhibit different motions. Therefore, instance-level segmentation is more suitable in this context, and we choose SAM2. Here, we conduct an experiment in which semantic labels are used to replace SAM2 outputs for generating RDPs. In particular, we utilize PSPNet [zhao2017pyramid] as a representative semantic segmentation model, trained on ADE20K [zhou2017scene] with its diverse and extensive category set. As shown in Tab. XI, this approach results in lower performance compared to our original implementation, supporting our claim.
IV-E Comparison in Low-light Conditions with VFI Baselines
Additionally, we compared our method against the original baselines in practical yet challenging scenarios, such as low-light environments. Tab. X presents the quantitative comparison results, while Fig. 10 provides visual comparisons. For these evaluations, we applied low-light degradations to the input frames from three datasets before testing. The comprehensive comparisons confirm the effectiveness of our framework in improving interpolation performance under low-light conditions.
This experiment also demonstrates that our method is robust to lighting variations (a common challenge in video interpolation) largely due to our use of SAM2 outputs, which are inherently resilient to lighting changes.
Figure 10: Visual comparison of UPR-Net [Alpher39] and UPR-Netours on SNU-FILM [Alpher43] whose data is degraded by the low-light.
IV-F User Study